# Stagewise Reinforcement Learning and the Geometry of the Regret Landscape

### v1

@@ Abstract
Singular learning theory characterizes Bayesian learning as an evolving tradeoff between accuracy and complexity, with transitions between qualitatively different solutions as sample size increases. We extend this theory to deep reinforcement learning, proving that the concentration of the generalized posterior over policies is governed by the local learning coefficient (LLC), an invariant of the geometry of the regret function. This theory predicts that Bayesian phase transitions in reinforcement learning should proceed from simple policies with high regret to complex policies with low regret. We verify this prediction empirically in a gridworld environment exhibiting stagewise policy development: phase transitions over SGD training manifest as “opposing staircases” where regret decreases sharply while the LLC increases. Notably, the LLC detects phase transitions even when estimated on a subset of states where the policies appear identical in terms of regret, suggesting it captures changes in the underlying algorithm rather than just performance.
@@ 1 Introduction
What determines the kinds of agents that deep reinforcement learning will produce? This question is both scientifically fundamental and consequential for real-world deployment: the inductive biases of RL training may underlie phenomena relevant to AI alignment ranging from goal misgeneralization to reward hacking (Anwar et al., 2024, §2.5). Yet stochastic gradient descent, the workhorse of deep RL, remains difficult to analyze directly. In this paper we study a more theoretically tractable learning process, Bayesian inference of parametrized policies, and prove a result with surprising implications: a more optimal policy is not always more optimal from a Bayesian perspective. Bayesian learning involves an evolving tradeoff between the regret of a policy and the complexity of its parametrization, as measured by an invariant of the geometry of the regret function known as the local learning coefficient (Watanabe, 2009; Lau et al., 2025). For finite data, a simpler policy with higher regret can outcompete a complex policy with lower regret. Understanding this tradeoff may illuminate how simplicity biases shape the agents that emerge from training.
Our approach is indirect: rather than analyzing SGD trajectories, we characterize which policies are favored by Bayesian inference and then ask how this relates to the outcomes of stochastic optimization. This strategy splits a hard problem into two parts. The first part – understanding Bayesian learning for deep RL – is the focus of this paper. The second part – understanding the relationship between concentration of the posterior and SGD dynamics – remains open and is an important area for future work.
To develop a Bayesian theory of deep RL we must confront two difficulties. First, the data in RL is not identically distributed: trajectories depend on the agent’s actions, so different agents observe different trajectories. Second, neural network policies are singular statistical models, meaning the Fisher information matrix is not invertible and standard asymptotic results (such as the Bernstein-von Mises theorem) do not apply. Singular learning theory (SLT; Watanabe, 2009) provides the mathematical tools to handle this second difficulty. Our contribution is to extend these tools to the RL setting, where both difficulties are present.
In supervised learning, SLT predicts that Bayesian learning should exhibit phase transitions: as sample size increases, the posterior can shift abruptly from concentrating in one region of parameter space to another (Watanabe, 2009; Chen et al., 2024b, §7.6). These transitions reflect the evolving tradeoff between accuracy and complexity encoded in Watanabe’s free energy formula. As we will show, the free energy formula also applies in RL, which leads to three predictions. First, the posterior over policies should exhibit phase transitions: sudden shifts from simpler, higher-regret policies to more complex, lower-regret policies as the agent incorporates more trajectory data. Second, these Bayesian transitions should have observable consequences for SGD training, manifesting as plateaus in return punctuated by rapid transitions. Third, such transitions should be accompanied by increases in the local learning coefficient, our complexity measure. The empirical results of this paper validate all three predictions.
We make the following contributions:
  • We define a generalized posterior over parameters for RL policies using importance sampling, connecting to the control-as-inference literature (Levine, 2018) and work of Bissiri et al. (2016) on belief updating (Section 2.4).
  • We prove that Watanabe’s free energy formula and WBIC theorem generalize to this setting under some assumptions on the environment (Section 2), establishing that the local learning coefficient (LLC) governs posterior concentration. This gives a precise sense in which the geometry of the regret function $G$ governs Bayesian learning for deep RL agents and allows us to make the aforementioned observation about more optimal policies (lower regret) not always being more optimal from a Bayesian perspective (higher posterior concentration), see Section 2.7.
@@
  • We use these results to derive a theoretical prediction: Bayesian phase transitions in RL should proceed from simple, high-regret policies to complex, low-regret policies (Section 2.7).
  • We verify this prediction empirically in a gridworld environment exhibiting stagewise learning, with “opposing staircases” of decreasing regret and increasing complexity (Figure 1).
  • We show that the LLC estimator detects phase transitions even when evaluated on a subset of states where the policies have identical regret, suggesting it captures algorithmic rather than purely behavioral changes (Section 3.4.4).
The remainder of the paper is organized as follows. Section 2 develops the theoretical foundations, defining the generalized posterior and proving our main theorem. Section 3 gives the empirical results: Section 3.2 presents the environment and the agent trained to solve it, Section 3.3 explains how we estimate the LLC and Section 3.4 presents our study of phase transitions. In Section 4 we review related work, and in Section 5 we return to discuss how these theoretical results allow us to think at a high-level about the kinds of agents that deep reinforcement learning will produce.
@@
2 Theory
2.1 Background
@@
By the Bayesian learning process we mean how the posterior distribution $p(w|D_{n})$ over parameters $w\in W$ given data $D_{n}$ (e.g. observed trajectories in an RL setting) evolves as $n$ increases. In supervised learning for “regular” models, where the data consists of $n$ random variables sampled independently and identically and the Fisher information matrix is invertible, the story is simple: the posterior converges as $n\rightarrow\infty$ to a normal distribution centered on the maximum likelihood estimator with variance controlled by the (inverse of the) Fisher information matrix by the Bernstein-von Mises theorem, see Hartigan, 1983.
However, the study of Bayesian learning for deep RL differs in two important ways from this simple story:
  • Non-identically distributed data: in RL the data from which an agent learns optimal behavior is typically a set of trajectories $o_{0},a_{0},r_{0},o_{1},a_{1},r_{1}\cdots$ consisting of actions $a$, observations $o$ and rewards $r$. This source is not identically distributed, as the actions of the agent shape the distribution of trajectories: for example in a video game, more experienced agents may gain access to new parts of the environment
@@
  • The model is singular rather than regular: when we parametrize our statistical model (in the case of RL, our policy) by a neural network the Fisher information matrix is not invertible and the model is not regular. This is the typical situation in deep learning Watanabe (2007); Wei et al. (2022b). Statistical models which are not regular are called singular and the appropriate mathematical framework for studying the Bayesian learning process for supervised learning is singular learning theory (SLT; Watanabe (2009)) which introduces ideas from algebraic geometry.
These two factors imply that a mathematical theory of parametric Bayesian learning which is general enough to treat deep RL must be nontrivial. Indeed, such a theory does not currently exist. The first contribution of this paper is to establish such a theory, in this section. The burden of additional mathematical sophistication has an upside: the Bayesian learning process is more interesting.
@@
In supervised learning it is known that the posterior for singular models, rather than converging to a normal distribution as in the regular case, can “jump” at certain critical sample sizes $n^{*}$ from concentrating in one region of parameter space for $n<n^{*}$ to concentrating in another region for $n>n^{*}$. This behavior resembles phase transitions in statistical physics (Watanabe, 2009, §7.6). In short, SLT for supervised learning predicts that the Bayesian learning process should involve phase transitions. This has been verified for Bayesian learning in small models (Chen et al., 2024b) and is conjecturally related to stagewise learning in SGD training in a range of systems including transformers (Hoogland et al., 2025). These phase transitions are explained by a theorem Watanabe (2013) that characterizes the optimal tradeoff between how well a model fits the data (accuracy) and its complexity, measured by a key quantity in SLT known as the local learning coefficient (LLC; Lau et al. (2025)). Stagewise learning has also been observed in deep reinforcement learning (Clift et al., 2020; McGrath et al., 2022).
2.2 Details
We will begin by discussing the foundational results on which our complexity measure for deep reinforcement learning models is based: a generalization of the theory of singular learning as developed by Watanabe Watanabe (2009; 2018) from distribution learning to a more general class of problems.
We will state our results for finite Markov decision problems in this setting, however the assumption of finiteness is not necessary; we treat a more general case in Appendix E. Our results also apply beyond the Markov decision problem setting to a general generalized Bayesian learning context that encompasses both reinforcement and supervised learning.
2.3 Setup
@@
It will be convenient to write the regret in the alternative form $\mathbb{E}_{\tau\sim q_{w}}(g(\tau))$ where $g(\tau)=R_{\mathrm{max}}-r(\tau)$.
@@ 2.4 The Generalized Posterior
    Reward vs probability: we follow the Kalman–Todorov duality Kalman (1960); Todorov (2008), which relates optimal control to optimal prediction under the following conceptual relation
@@
    This has been studied more recently under the name “control as inference” (Levine, 2018).
@@
Given a prior $\phi(w)$ over parameters for policies, to complete a Bayesian description of reinforcement learning it only remains to define how this distribution is updated given observations $D_{n}$. That is, we need to define a generalized posterior $p(w|D_{n})$. Since $g(\tau)$ is our cost function and the $\tau_{i}$ are independent, (1) gives in the on-policy case (where $w_{i}=w_{0}$ for $1\leq i\leq n$)
@@
where $G_{n}(w_{0})=\frac{1}{n}\sum_{i=1}^{n}g(\tau_{i})$. However we can’t use Bayes rule to obtain $p(w|D_{n})$, because we can only define $p(D_{n}|w)$ this way for $w=w_{0}$. To proceed note that $G_{n}(w_{0})$ is a Monte Carlo estimator for
@@
where the normalizing constant $Z$ is chosen so that this is a probability distribution. Note that this reduces to the on-policy calculation above when $w_{i}=w_{0}$ for all $i$ and $\beta=1$. We will also need to introduce a tempered version of this posterior, and notation for the probability of an open set according to the posterior:
@@
For more discussion of the generalized posterior see Section E.2.
2.5 Main Theoretical Results
@@
Our results generalize Watanabe (2018, Section 6.3) and Watanabe (2013, Theorem 4) to RL.
@@
Our main theorem will require imposing the following condition on a Markov decision problem. If $w\in W$ is an optimal parameter, we will assume that $g(\tau)=0$ almost-always for the distribution $q_{w}$. In other words optimal policies almost always receive optimal return. This is true, for example, if the transition function is deterministic.
@@
Consider a Markov decision problem satisfying Assumption 2.2. Then the generalized posterior obeys the following conditions:
@@
    where $\lambda(U),m(U)$ are the learning coefficient and multiplicity of $G$ on the set $U$  [^2] The quantities $\lambda(U)$ and $m(U)$ are referred to respectively by algebraic geometers as the real log canonical threshold and the real log canonical multiplicity of the function $G$ restricted to the set $U$..
@@
Proof.
See Appendix E, specifically E.2.4 and E.2.5. ∎
@@ 2.6 The Local Learning Coefficient
Given a point $w^{*}$ in parameter space with open neighborhood $U$, according to Theorem 2.3 we may identify $\lambda(U)$ as
$\lambda(U)=n\beta\left[\mathbb{E}_{n,U}^{\beta}[G_{n}(w)]-G_{n}({w^{*}})\right],$ (7)
@@
More geometrically one can equivalently realize the local learning coefficient in terms of the asymptotic growth rate of basin volumes around a local minimum (see (Lau et al., 2025, §3.1)).
Theorem 2.3 explains how the local learning coefficient controls model selection during the learning process. Taking $\beta=1/\log n$ and writing $F_{n}=F_{n,1}$, Parts (1) and (2) of Theorem 2.3 together imply (see Remark E.44) the free energy formula for a local minimum $w_{0}$ of $G$

### v2

@@ Abstract
Singular learning theory characterizes Bayesian learning as an evolving tradeoff between accuracy and complexity, with transitions between qualitatively different solutions as sample size increases. We extend this theory to reinforcement learning, proving that the concentration of a generalized posterior over policies is governed by the local learning coefficient (LLC), an invariant of the geometry of the regret function. This theory predicts that deep reinforcement learning with SGD should proceed from simple policies with high regret to complex policies with low regret. We verify this prediction empirically in a gridworld environment exhibiting stagewise policy development: phase transitions over training manifest as “opposing staircases” where regret decreases sharply while the LLC increases.
@@ 1 Introduction
What principles determine the kinds of agents that deep reinforcement learning produces? This question grows increasingly salient as these principles shape the agents that play ever more integral roles in our society with each passing year. This question also presents a fundamental challenge for deep learning science, since its answering requires understanding the inductive biases of the kinds of extremely complex high-dimensional stochastic optimization processes that comprise neural network training.
In light of the formidable nature of this challenge, we propose the following approach to an empirical understanding of the inductive biases of deep RL algorithms: First, we should theoretically understand the inductive biases of an idealized learning process chosen to capture core elements of deep RL, while still being analytically tractable. Second, we should derive predictions about the kinds of phenomena we would see in deep RL if it were driven by similar principles. Finally, we should experimentally test these predictions in order to validate the idealized learning process as a partial model of the real deep RL process.
A similar approach has recently proven fruitful in the science of deep supervised learning (e.g, Chen et al., 2023; Hoogland et al., 2025; Carroll et al., 2025). These works use singular Bayesian inference as an idealized learning process, applying Watanabe’s free energy formula for Bayesian inference in singular statistical models (Watanabe, 2009; 2018) to frame learning as an evolving tradeoff between accuracy (as measured by loss) and complexity (as measured by the local learning coefficient, LLC, a geometric invariant of the loss landscape; Lau et al., 2025). With this understanding, they predict and experimentally verify the existence of various stagewise learning phenomena in deep learning and their connection to loss landscape geometry.
In this paper, we extend this approach to deep RL, offering the following contributions:
  1. 1.
    In Section 3, we develop our idealization of the deep RL process, namely a generalization of singular Bayesian inference that accounts for the role played by rewards in RL using Kalman–Todorov duality (Kalman, 1960; Todorov, 2008; Levine, 2018), as well as accounting for potential non-stationarity in experience data using an importance sampling objective.
  2. 2.
    In Section 4, we outline a generalization of Watanabe’s free energy formula and Bayesian information criterion for singular statistical models, extending the main results of singular learning theory to our generalized inference setting. The proof follows Watanabe (2009; 2013; 2018), with the main extensions being the application of Lyapounov’s central limit theorem and the identification of a new fundamental condition. Appendix D gives proofs of these results in a unified setting.
  3. 3.
@@
    In Section 5, we leverage these new results to derive qualitative predictions about deep RL. In particular, we predict that deep RL should proceed stagewise from simple, high-regret policies to complex, low-regret policies, with policy complexity measured by the LLC.
  4. 4.
    In Section 6, we validate these predictions in deep reinforcement learning experiments in a simple gridworld environment. Figure 1 previews the “opposing staircases” phenomenon, where we see the policy undergoes stagewise learning with decreasing regret and increasing complexity as measured by LLC.
Our results demonstrate that deep reinforcement learning shares a fundamental property of Bayesian inference: the agents produced by learning are not necessarily optimal according to the return/regret. Rather, deep reinforcement learning agents exhibit stagewise learning governed by an evolving tradeoff between performance and simplicity. This learning dynamic has important consequences for practical applications of deep reinforcement learning in current and future agentic AI systems. We discuss these implications in Section 7.1. Our results also underscore the LLC as a principled model complexity measure for RL, we survey related complexity measures in Section 7.2. Finally, our results provide a stepping stone for further theoretical analysis of the dynamics of deep RL, and invite the development of SLT-based interpretability and training tools, which we discuss in Section 7.3.
@@
2 Background
In this section, we provide a brief primer on singular learning theory (SLT) for supervised learning, and introduce preliminary definitions for discussing reinforcement learning (RL).
2.1 Bayesian Deep Supervised Learning and Singular Learning Theory
When we refer to Bayesian inference as a learning process, we mean how the posterior distribution $p(w|D_{n})$ over parameters $w\in W$ given data $D_{n}$ evolves as $n$ increases.
@@ 2.1 Bayesian Deep Supervised Learning and Singular Learning Theory
In supervised learning for regular statistical models, where the data consists of $n$ random variables sampled independently and identically and the Fisher information matrix is invertible, the story is well-known and simple: as $n\rightarrow\infty$, the posterior converges to a normal distribution centered on the maximum likelihood estimator with variance controlled by the (inverse of the) Fisher information matrix by the Bernstein-von Mises theorem, see Hartigan, 1983.
@@
However, when we parametrize our statistical model by a neural network the Fisher information matrix is not invertible and the model is not regular. This is the typical situation in deep learning (Hagiwara et al., 1993; Watanabe, 2007; Wei et al., 2022b). Statistical models which are not regular are called singular and the appropriate mathematical framework for studying the Bayesian learning process for supervised learning is singular learning theory (SLT; Watanabe, 2009), which leverages algebraic geometry to describe Bayesian learning dynamics in terms of the nature of singularities in the negative log likelihood.
In particular, SLT gives an asymptotic expansion of the local free energy, known as Watanabe’s free energy formula, which we informally describe now. Let $u\in W$ be a local minimum of the expected negative log likelihood, and let $U$ be a ball around $u$ in which $u$ is a global minimum. Then, under certain technical conditions on the statistical model, we have an asymptotic expansion in $n$ Watanabe, 2018, Thm. 11; Lau et al., 2025 of the free energy associated to this neighborhood:
$\log\int_{U}\exp(-nL_{n}(w))\phi(w){\rm d}w=nL_{n}(u)n+\lambda(U)\log n-(m(U)-1)\log\log n+O_{\mathrm{P}}(1)$
where $L_{n}(u)$ is the negative log likelihood of $D_{n}$ at $u$ and $\lambda(U)$ and $m(U)$ are coefficients derived from the geometry of the log likelihood around $u$. The coefficient $\lambda(U)$ is also known as the local learning coefficient (LLC; Lau et al., 2025). The lower-order terms include, for example, a contribution from the prior $\phi$.
@@
As we will discuss in later sections, this formula implies that the learning dynamics of Bayesian inference with singular models (including Bayesian deep supervised learning) are much richer than in the regular case. For example, rather than converging to a normal distribution as in the regular case, in singular models the posterior can “jump” at certain critical sample sizes $n^{*}$ from concentrating in one region of parameter space for $n<n^{*}$ to concentrating in another region for $n>n^{*}$. This behavior resembles phase transitions in statistical physics (Watanabe, 2009, §7.6), and these phase transitions are characterized by an evolving tradeoff between how well a model fits the data as measured by the log likelihood, and its complexity as measured by the LLC Watanabe, 2009, §7.6; Watanabe, 2013; Chen et al., 2023. This theory has been verified for Bayesian learning in small models (Chen et al., 2023) and is conjecturally related to stagewise learning in SGD training in a range of systems including transformers (Hoogland et al., 2025; Carroll et al., 2025).
So much for supervised learning. The main aim of this paper is to extend the above story to a more general setting that captures the core features of deep RL, including (1) the essential role of singularities, (2) the use of rewards rather than labels, and (3) the non-stationarity of experience data.
2.2 Reinforcement Learning Preliminaries
We now introduce the preliminary definitions we require for describing our results in the main text. Note that in Appendix D we treat a more general setting, relaxing some of the restrictions introduced here for brevity (such as finiteness of the Markov decision process).
@@ 2.2 Reinforcement Learning Preliminaries
Minimizing regret is equivalent to maximizing return. For simplicity, we frame our theory around minimizing regret. Moreover, it will be convenient to write the regret in the alternative form $\mathbb{E}_{\tau\sim q_{w}}(g(\tau))$ where $g(\tau)=R_{\mathrm{max}}-r(\tau)$.
3 Bayesian Deep Reinforcement Learning
In this section, we develop our idealized learning process, a generalization of Bayesian inference that accounts for the role of rewards and the non-stationarity of experience data in reinforcement learning. Section 3.1 culminates in Definition 3.1, giving the form of the generalized Bayesian posterior (or “Gibbs posterior”). Section 3.2 discusses the relationship between this idealized learning process and SGD-based deep RL.
@@ 3.1 The Generalized Posterior
    Reward versus probability: we follow the Kalman–Todorov duality (Kalman, 1960; Todorov, 2008), which relates optimal control to optimal prediction under the following conceptual relation
@@
    This principle has been studied more recently under the name “control as inference” (Levine, 2018).
@@
Given a prior $\phi(w)$ over parameters for policies, to complete a Bayesian description of reinforcement learning it only remains to define how this distribution is updated given observations $D_{n}$. That is, we need to define a generalized posterior $p(w|D_{n})$. Since $g(\tau)$ is our cost function and the $\tau_{i}$ are independent, (1) gives a likelihood in the on-policy case (where $w_{i}=w_{0}$ for $1\leq i\leq n$)
@@
where $G_{n}(w_{0})=\frac{1}{n}\sum_{i=1}^{n}g(\tau_{i})$. However we can’t use Bayes rule to obtain a posterior $p(w|D_{n})$, because we can only define a likelihood $p(D_{n}|w)$ this way for $w=w_{0}$. To proceed note that $G_{n}(w_{0})$ is a Monte Carlo estimator for
@@
where the normalizing constant $Z$ is chosen so that this is a probability distribution. Note that this reduces to the on-policy calculation above when $w_{i}=w_{0}$ for all $i$. We will also need to introduce a tempered version of this posterior, and notation for the probability of an open set according to the posterior:
@@
For more discussion of the generalized posterior see Section D.2.
3.2 Relationship to Deep Reinforcement Learning
$\bigl\{(w^{(k)},\tau^{(k)}_{j})\;:\;k\leq K,\;1\leq j\leq B_{k}\bigr\},$
i.e., it is a union of samples drawn under earlier iterates of the same SGD process. [^2] In distributed actor–learner architectures one should interpret $w^{(k)}$ as the parameters on the actor at the moment of rollout, which are typically a lagged copy of the learner parameters. Thus, even though our theory allows an arbitrary sequence $(w_{i})_{i=1}^{n}$, the “canonical” source of such a sequence in practice is simply the learning algorithm itself, via checkpoints or lagged actors.
which matches the on-policy training rule used in this paper (cf., 15).
and replacing $g(\tau_{i})$ by its reward-to-go form yields the trajectory-sampled analogue of (11) (up to the overall sign corresponding to maximizing return versus minimizing regret). Thus, stochastic gradient descent on $G_{n}$ corresponds to an off-policy REINFORCE-style update, while the generalized posterior is the Gibbs distribution associated to the same empirical objective.
Beyond this connection at the level of objectives, it is an open challenge to characterize the precise relationship between concentration of the generalized posterior and SGD dynamics. This represents an important next step for future theoretical work (see Section 7.3). However, for present purposes, we can immediately analyze the learning dynamics of the idealized case, treat this as a source of empirical predictions about SGD training, and then validate these predictions with deep RL experiments. We now turn to this analysis.
4 The Free Energy Formula
In this section, we outline our generalization of Watanabe’s free energy formula and Bayesian information criterion for singular statistical models, Watanabe, 2018, §6.3; Watanabe, 2013, Thm. 4. In particular, Theorem 4.2 extends these two results to the generalized inference setting developed in Section 3.
@@ 4 The Free Energy Formula
Our main theorem requires imposing the following fundamental condition on the Markov decision problem. If $w\in W$ is an optimal parameter, we assume that $g(\tau)=0$ almost-always for the distribution $q_{w}$. In other words optimal policies almost always receive optimal return. This is true, for example, if the transition function is deterministic.
@@
Consider a Markov decision problem satisfying 4.1. Then the generalized posterior obeys the following conditions:
@@
    where $\lambda(U)$, $m(U)$ are the learning coefficient and multiplicity of $G$ on the set UU. [^3] The quantities $\lambda(U)$ and $m(U)$ are referred to respectively by algebraic geometers as the real log canonical threshold and the real log canonical multiplicity of the function $G$ restricted to the set $U$.
@@
We defer proof to Appendix D, specifically D.2.4 and D.2.5. There, we offer a proof following and extending the arguments from Watanabe (2009; 2013; 2018). There are many details, but the main extensions are in the application of Lyapounov’s central limit theorem to address the use of independently but non-identically distributed data, and the identification of 4.1 as the only obstruction to lifting good local coordinates from $G$ to $G_{n}$ (this condition plays the role of the relative finite variance condition in Watanabe, 2018).
In fact, the treatment in Appendix D is more general than the statement above in several ways. First, we relax the finite Markov decision problem restriction. Second, we work in a unified setting that simultaneously encompasses both reinforcement learning and supervised learning. Third, we identify a more general fundamental condition from which we recover 4.1 in the RL setting.
5 Implications of the Free Energy Formula
In this section, we leverage Theorem 4.2 to derive qualitative predictions about deep reinforcement learning. We first elaborate on the local learning coefficient (Section 5.1) and the free energy formula (Section 5.2) and how they combine to describe stagewise development in the generalized posterior (Section 5.3). Then, in Section 5.4, we interpret this theory to form empirically testable predictions about how deep reinforcement learning should proceed stagewise from simple, high-regret policies to complex, low-regret policies, with policy complexity measured by the LLC.
@@ 5.1 The Local Learning Coefficient
In the Bayesian setting, the LLC provides a rigorous formalization of the notion of the complexity of a particular parameter. Given a point $w^{*}$ in parameter space with open neighborhood $U$, according to Theorem 4.2 we may identify $\lambda(U)$ to leading order in $n$ with the following expression:
$\lambda(U)\sim n\beta\left[\mathbb{E}_{n,U}^{\beta}[G_{n}(w)]-G_{n}({w^{*}})\right],$ (13)
@@
More geometrically one can equivalently realize the local learning coefficient in terms of the asymptotic growth rate of basin volumes around a local minimum (see Lau et al., 2025, §3.1).
5.2 The Generalized Free Energy Formula
Theorem 4.2 explains how the local learning coefficient controls model selection during the learning process. Taking $\beta=1/\log n$ and writing $F_{n}=F_{n,1}$, Parts (1) and (2) of Theorem 4.2 together imply (see Remark D.44) the free energy formula for a local minimum $w_{0}$ of $G$